### v2

@@ Discussion
We demonstrated GHz-fast EO phase shift and modulation in FN-LC. In contrast to EO polymers, we established that the material does not require any electro-thermal poling processes for alignment. Indeed, the material can be oriented during regular modulator RF operation. An FNS waveguide infiltrated with this material is proposed to reduce the number of lithography steps compared to slot waveguides while benefiting from improved LMI compared to non-slot ones. Using FNS phase shifters as small as 500 in a push-pull MZM, we achieved a DC modulation efficiency of 0.25 Vmm, an optical ER of 26 dB, and an IL of 2.1 dB. The propagation loss of dBmm-1 is still in the order of some of slot waveguides [62-63]. We discovered that the material absorption contributes 0.5 dB of the total IL and is very similar to the case of an EO polymer. The majority of the loss is attributed to the non-optimized FNS waveguide. Reducing fabrication imperfections, including less core sidewall roughness and better infiltration of FN-LC into the two spaces between the core and fingers, can further improve IL. A significant phase shift efficiency in the order of 88.9 was estimated, which we attribute to the material and the FNS waveguide employed in this work. Most importantly, we discovered a high-speed phase shift mechanism with an EO bandwidth > 4.18 GHz that outperforms the stronger but slower birefringence effect found in classic paraelectric-nematic liquid crystals. The EO bandwidth of our FN-LC-covered modulator is limited by the experiment setup and device fabrication, not the suggested architecture or the material, both of which are easy to fix but are out of the scope of this work.
Also, an analytical approach was introduced to estimate the Pockels coefficient of the material, where an AC modulation efficiency of 25.7 Vmm associated with a modest 24 pmV-1 at MHz. As a sufficiently reliable and stable material for practical applications that can scale, the comparatively low modulation efficiency at higher speeds for this first-ever prototype is still a breakthrough and a harbinger of further improvements. Other waveguide embodiments featuring stronger LMI can be coated with FN-LC to achieve high AC modulation efficiencies at the cost of more involved fabrication processes.
@@
The coexistence of a high-speed and effective but slow phase shift mechanism in FN-LC could pave the way for poling-free silicon organic PICs that have similar device metrics to those of EO polymers but require fewer strict steps to be made. The non-thermal nature of the alignment step allows for the selective alignment of phase shifters without affecting adjacent devices, which is one of the main benefits of this yet-to-be-thoroughly validated result. Future research will explore the performance of our proposed poling-free Si-organic modulators based on FN-LC for high-speed data transmission, switching, or transduction at cryogenic temperatures. One hypothesis is to leverage the potentially better material’s elasticity than EO polymers. It could be advantageous to relax the tension caused by different thermal expansions in such a SiP chip without requiring surface passivation or material alterations. Future research will examine the material’s EO properties (i.e., , and ) at cryogenic temperatures.
@@ Methods
Device fabrication: The modulators were patterned on a standard 220 nm silicon on an insulator wafer with a buried oxide layer of 3.5 . An electron beam lithography tool (JBX-8100FS) was used to design the waveguides via positive tone resist (ZEP 520A-7) [56-57]. After that, the devices were dry-etched with an ICP plasma etcher tool (Oxford COBRA). Using maskless lithography (Heidelberg MLA-150), metal electrodes (100 nm gold atop 5 nm titanium as an adhesion layer) were deposited 4.85 m distant from the core waveguide. A passive optical measurement was performed for reference after completing the nanofabrication stages of photonic-only structures and before metallization.
@@
After metallization, both chips were cleaned gently using acetone, IPA, DI water, and a N2 gun, followed by a 30 sec oxygen plasma cleaning for surface passivation just before coating. For the OEO cladding, we used PM-158 (a proprietary FN-LC material optimized for enhanced obtained from Polaris Electro-Optics, Inc.). According to our pre-coating viscosity experiments, the cleaned chips benefitted from being warmed up (70C) to reduce the viscosity of FN-LC during application to the chip. A glass capillary tube was used to apply enough FN-LC to the desired areas. Finally, the chips were cooled to 25∘C before characterization.
FN-LC alignment and DC characterization: While no significant relationship was found between the timing of the alignment stage and the time passed after coating FN-LC, the alignment step was performed on the same day. After coating, the produced samples were mounted on an active EO probe station with temperature control. A source meter unit (Keithley 2400) was employed to apply step voltage and monitor polarization-induced current. This step should be done before AC characterization utilizing the bias Tee shown in Fig. 4(a). As illustrated in Fig. 3(b), we found out a poling field of should be enough for complete alignment, which is at least 4 times smaller than most of existing EO polymers [53,64-65]. The light from a tunable CW laser (Agilent 8164A/81682A) is coupled into grating couplers using an 8-degree polished angle FA. In this phase, we tune the laser wavelength to the highest peak achievable by overlapping the MZM spectrum with that of the GC or PWB. An InGaAs power sensor (Agilent 81635A) detects light. Realtime measurement of optical loss during FN-LC alignment is straightforward and advantageous over EO polymer poling, which requires more advanced control due to elevated temperatures. It can serve as an additional indication of successful alignment. The components of phase shift efficiency (i.e., and ) were extracted using the associated shifts in the vertical () and horizontal () directions as a function of applied voltage.
@@
AC characterization: The same SMU is used as a DC source with a set voltage precision of 0.02 + 24 mV, which is 20 times greater than , thus allowing precise adjustment of MZM bias despite the efficient DC detuning mechanism in FN-LC. An 18 GHz vector network analyzer (Keysigth FieldFox VNA) was added for AC characterization, with port 1 amplified using a microwave amplifier (SHF-s807). It is mixed with a proper amount of DC voltage for optimum MZM performance and simultaneously for FN-LC alignment. A high-speed photodetector (Thorlabs RXM40AF) was employed at the output, along with low-noise amplifiers (picosecond 5828). The VNA data was utilized to evaluate the device S-parameters. The devices were aged at a regulated temperature (25∘C) for 150 hours with no humidity control or hermetic sealing applied, with light ( = 8 dBm) and an RF electrical signal ( = 1.59 V) applied.
@@
The material EO coefficient of FN-LC (i.e., ) was post-processed using the analytical formula [52] (Supplementary Note 1):
@@
where is the calibrated gain of the modulator link, is the total optical loss, is the optical output of the MZM, () is the optical (electrical) input (output) of the detector, is the effective index of the FNS waveguide covered with FN-LC, and is the MZM drive voltage, respectively. Here, and correspond to the detector and the MZM sensitivities. The latter is related to and, within good accuracy, to by:
@@
In Equation 4, n is the field-dependent effective mode index of FNS waveguide, A and V are the cross-section and volume of the LMI region, Z0 is the free space impedance, and is the magnetic field of the optical mode, respectively.
Starting with an approximation (e.g., [67,68]), an electric field simulator is utilized to generate field variable refractive index, which is iteratively called by an optical mode solver (COMSOL Multiphysics). After finite iterations, the solution for the overlap integral should converge to a unique number for each architecture. Both electrical and optical modal field solvers are in 3D to consider the effect of the periodic structure of the FNS waveguide along the light propagation. The semi-double slot waveguide structure of the FNS waveguide can offer compared to of a slot waveguide (=100 nm) (Supplementary Note 2).
However, the characteristic figure of merit of a phase shifter () is . We have also compared a hypothetical nonslotted waveguide with a 300-nm core width surrounded by 2.1 trenches. Our simulations revealed 0.26, 0.27, 0.15 for the proposed FNS waveguide, slotted and nonslotted structures, respectively. If doped properly, one can assume = 200 nm, i.e., 21 times smaller than that of a non-slotted structure. Our 3D point-by-point evaluations of agreed well with this approximation. On the other hand, we have recorded a total propagation loss of dB/mm for the FNS waveguide, versus dB/mm for the above-described, hypothetical nonslotetd structures within the best etching and coating quality we could achieve. Using these three ingredients, one can conclude an improvement of 8.7 times compared to a non-finger-loaded waveguide. The metrics are estimated to be still at least twice higher for the aforementioned slot waveguide thanks to smaller , at the cost of the described challenges.
@@
The algorithms used for this study are standard and are outlined in “Methods”. Code scripts can be provided by the corresponding authors upon reasonable request.
@@ Acknowledgements
The Natural Sciences and Engineering Research Council of Canada (NSERC), the SiEPICfab consortium, the B.C. Knowledge Development Fund (BCKDF), the Canada Foundation for Innovation (CFI), MITACS, and the Schmidt Sciences funded this research.
@@
This work was conceptualized by I.T., L.C., S.S. C.P., J.S., and N.J. contributed materials and technical information. At the same time, S.J., D.W., and I.T. built the organic Si modulator described in this paper. I.T. conducted measurements while N.J. and O.E. helped with the modulator measurement setup. The manuscript was written by I.T., L.C., and S.S., with assistance from all authors.
@@
e and f in the y-slice perspective (y = 110 nm). Compared to a non-slotted waveguide, the light-matter interactions (LMI)-engineered FNS waveguide acts like a double-slot waveguide to offer 21 times improvement in the as a result of decreased effective electrode distance () as well as 1.75 times higher . () drops to before it overlaps with the encompassing fingers. g A zoomed-in view of indicating that both and are evaluated at each coordinate (x,y,z) in the space, where light interacts with the FN-LC. The marked point proved to be a good reference when estimating (see Methods). h evaluated at 4.18 GHz indicates the RF loss originated by low doping of the FNS waveguide (cm-3).
c A microscopic image of the MZM along with the dipolar order of FN-LC in a section of its finger-loaded non-slotted (FNS) waveguide in the cases of and (other possibilities for dipolar order could exist, since we don’t know what the ground state is with this topology). d Optical output spectrum (normalized) versus applied voltage exhibiting changes in the EOPS complex refractive index ( where for an ideal EOPS). The data is summarized in Table LABEL:FOMtable.
b MZM schematic showing two 500 m-long EOPS and low-loss strip-to-finger-loaded mode converters equipped with GC and PWB optical I/O channels. The alignment has been performed unipolarly, as evidenced by the green arrows. However, as the dark blue arrows show, the RF field is applied bipolarly to achieve push-pull performance. The proposed EOPS made of FN-LC incorporates both slow and fast phase shift mechanisms compared to the conventional approach, shown in c, in which an inefficient thermo-optic effect is required in conjunction with a fast EOPS for advanced modulation schemes. d Measured reveals the existence of a high-speed phase shift mechanism with an EO bandwidth of GHz. Post-processed AC values of based on extracted as a function of frequency are also shown. e A false-colored scanning electron microscope (SEM) image of the Light-matter interaction (LMI)-engineered finger-loaded non-slotted (FNS) waveguide with a zoomed-in view of its mode converter.

### v3

@@ Discussion
We demonstrated GHz-fast EO phase shift and modulation in FN-LC. In contrast to EO polymers, we established that the material does not require any electro-thermal poling processes for alignment. Indeed, the material can be oriented during regular modulator RF operation. An FLS waveguide infiltrated with this material is proposed to reduce the number of lithography steps compared to slot waveguides while benefiting from improved LMI compared to non-slot ones. Using FLS phase shifters as small as 500 in a push-pull MZM, we achieved a DC modulation efficiency of 0.25 Vmm, an optical ER of 26 dB, and an IL of 2.6 dB. The propagation loss of is similar to some slot waveguides [60-61]. We discovered that the material absorption contributes 0.5 dB of the total IL and is very similar to the case of an EO polymer. The majority of the loss is attributed to the non-optimized FLS waveguide. Reducing fabrication imperfections, including less core sidewall roughness and better infiltration of FN-LC into the two spaces between the core and fingers, can further improve IL.
A significant phase shift efficiency in the order of 88.9 was estimated, which we attribute to the material and the FLS waveguide employed in this work. Most importantly, we discovered a high-speed phase shift mechanism of at least 4.18 GHz that outperforms the more substantial but slower birefringence effect in classic paraelectric-nematic liquid crystals. The EO bandwidth of our FN-LC-covered modulator is limited by the experiment setup and device fabrication, not the suggested architecture or the material, both of which are easy to fix but are out of the scope of this work. We established the existence of a Pockels-based linear electro-optic effect in FN-LC materials.
Also, an analytical approach was introduced to estimate the Pockels coefficient of the material, where an AC modulation efficiency of 25.7 Vmm associated with a modest 24 pmV-1 at MHz. The difference between at DC and AC originates from the phase shift mechanism, which is dominant at each regime. At DC, the molecular orientation causes a significant modulation sensitivity (i.e., ). However, the faster Pockels effect dominates at higher frequencies, influenced by the modest reported above. Doping silicon waveguides can also lower , resulting in a higher voltage drop ratio over the LMI region (i.e., , where f is the operating frequency, and represent the resistance of the finger and capacitance of the gap between the core and finger, respectively). The effect is less significant unless at low frequencies, at which , where .
@@
As a sufficiently reliable and stable material for practical applications that can scale, the comparatively low modulation efficiency at higher speeds for this first-ever prototype is still a breakthrough and a harbinger of further improvements. Other waveguide embodiments featuring stronger LMI can be coated with FN-LC to achieve high AC modulation efficiencies at the cost of more involved fabrication processes. The coexistence of a high-speed and effective but slow phase shift mechanism in FN-LC could pave the way for poling-free silicon organic PICs that have similar device metrics to those of EO polymers but require fewer strict steps to be made. The non-thermal nature of the alignment step (i.e., poling-free) allows for the selective alignment of phase shifters without affecting adjacent devices, which is one of the main benefits of this yet-to-be-thoroughly validated result. Future research will explore the performance of our proposed poling-free Si-organic modulators based on FN-LC for high-speed data transmission, switching, or transduction at cryogenic temperatures. The material will likely lose its beneficial DC phase shift as temperature decreases. Besides, alignment should be maintained during the cool-down process. Nonetheless, one may leverage the potentially better material’s elasticity than EO polymers. It could be advantageous to relax the tension caused by different thermal expansions in such a SiP chip without requiring surface passivation or material alterations. Future research will examine the material’s EO properties (i.e., , and ) at cryogenic temperatures.
@@ Methods
Device fabrication: The modulators were patterned on a standard 220 nm silicon on an insulator wafer with a buried oxide layer of 3.5 . The pattern was first corrected for proximity effects using GenIsys Beamer software, applying both short- and long-range proximity effect corrections. The silicon-on-insulator (SOI) sample was cleaned using a standard solvent cleaning process with acetone and isopropyl alcohol (IPA), followed by nitrogen drying and a 30-second oxygen plasma treatment (20 sccm O2 flow). The sample was then baked at 180∘C for dehydration. Zep 520A-7, a positive-tone electron beam lithography (EBL) resist, was spin-coated at 2000 rpm with 1000 rpm/s for 35 seconds to achieve a resist thickness of approximately 260 nm. The sample was subsequently baked at 180∘C to remove any residual solvent. [54-55].
Resist exposure was carried out using a 100 keV EBL system (JEOL JBX-8100FS, JEOL) with a dosage of 213 , a beam current of 1 nA, and a shot pitch of 4 nm. Development was performed in N-Amyl Acetate (ZED-N50, Zeon Specialty Materials Inc.) for 1 minute. The pattern was then transferred into the silicon device layer using an inductively coupled plasma reactive ion etching (ICP-RIE) process (Oxford COBRA). Etching parameters included a chamber pressure of 10 mTorr, a temperature of 15∘C, an ICP power of 600 W, a bias power of 30 W, and a bias voltage of 220 V. The gas mixture consisted of 30 sccm of C4F8 and 25 sccm of SF6, with a total etching duration of 3.5 minutes. Residual Zep resist was removed by exposing the sample to deep ultraviolet light for 15 minutes, followed by immersion in a heated PG remover bath for 15 minutes. The sample was then rinsed in IPA and dried with nitrogen.
For metallization, the sample was treated again with a 30-second oxygen plasma (20 sccm O2 flow) and baked at 110∘C for dehydration. AZ514E photoresist was used in image reversal mode for the lift-off process. Spin coating was performed at 4000 rpm with an acceleration of 1000 rpm/s for 40 seconds. After a post-bake at 110∘C for 1 minute, the sample was exposed using a maskless aligner (MLA-150, Heidelberg Instruments) with a 365 nm wavelength and an exposure dose of 40 mJ/cm2. The sample underwent a second bake at 90∘C for 1 minute, followed by a flood exposure. Development was carried out using a dilute TMAH solution (AZ MIF 300 Developer, MicroChemicals) for 1 minute, followed by a 1-minute rinse in deionized water. Metal electrodes (100 nm gold atop 5 nm titanium as an adhesion layer) were deposited 4.85 m distant from the core waveguide. A passive optical measurement was performed after completing the nanofabrication stages of photonic-only structures and after metallization. No evidence of EO activity was observed before coating FN-LC, which ensures no detectable contribution from free carrier or plasma dispersion effects.
@@
After metallization, the chip is cleaned gently using acetone/IPA and an additional step of DI water for removing any water-soluble residues, dried with N2 gun, and surface-passivated by a step of oxygen plasma treatment for 30 sec just before coating. For the OEO cladding, we used PM-158 (a proprietary FN-LC material optimized for enhanced obtained from Polaris Electro-Optics, Inc.). According to our pre-coating viscosity experiments, the cleaned chip benefitted from being warmed up (70C) to reduce the viscosity of FN-LC during application to the chip. A glass capillary tube was used to apply enough FN-LC to the desired areas. Finally, the chip is cooled to 25∘C before characterization.
@@
AC characterization: The same SMU is used as a DC source with a set voltage precision of 0.02% + 24 mV, which is 20 times greater than , thus allowing precise adjustment of MZM bias despite the efficient DC detuning mechanism in FN-LC. An 18 GHz vector network analyzer (Keysight FieldFox VNA) was added for AC characterization, with port 1 amplified using a microwave amplifier (SHF-s807). It is mixed with a proper amount of DC voltage for optimum MZM performance and simultaneously for FN-LC alignment. A high-speed photodetector (Thorlabs RXM40AF) was employed at the output, along with low-noise amplifiers (picosecond 5828). The VNA data was utilized to evaluate the device S-parameters. The devices were aged at a regulated temperature (25∘C) for 150 hours with no humidity control or hermetic sealing applied, with light ( = 8 dBm) and an RF electrical signal ( = 1.59 V) applied. In Fig. 4(e), an arbitrary function generator drives the amplifier, and the data captured by a digital storage oscilloscope (SDS6204A by SIGLENT) performs fast Fourier transformation.
@@
The material EO coefficient of FN-LC (i.e., ) was post-processed using the analytical formula [52] (Supplementary Note 1)
@@
where is the calibrated gain of the modulator link, is the total optical loss, is the optical output of the MZM, () is the optical (electrical) input (output) of the detector, is the effective index of the FLS waveguide covered with FN-LC, and is the MZM drive voltage, respectively. Here, and correspond to the detector and the modulation sensitivities. The latter is related to and, within good accuracy, to by:
@@
In Equation 4, n is the field-dependent effective mode index of FLS waveguide, A and V are the cross-section and volume of the LMI region, Z0 is the free space impedance, and is the magnetic field of the optical mode, respectively. As a first approximation, one can assume [65,66]. To obtain a more precise value for , an electrical field simulator calculates the local values of , the results of which are called by an optical mode solver (COMSOL Multiphysics) to incorporate the field dependency of . We assumed FN-LC has similar permittivities at the optical and microwave frequencies in analogy to EO polymers [70]. Both electrical and optical modal field solvers are in 3D to consider the effect of the periodic structure of the FLS waveguide along the light propagation. The semi-double slot waveguide structure of the FLS waveguide can offer compared to of a slot waveguide (=100 nm) (Supplementary Note 2).
However, the characteristic figure of merit of a phase shifter () is . We have also compared a hypothetical non-slotted waveguide with a 300-nm core width surrounded by 2.1 trenches. Our simulations revealed 0.26, 0.27, 0.15 for the proposed FLS waveguide, slotted and non-slotted structures, respectively. If doped properly, one can assume = 200 nm, i.e., 21 times smaller than that of a non-slotted structure. Our 3D point-by-point evaluations of agreed well with this approximation. On the other hand, we have recorded a total propagation loss of dB/mm for the FLS waveguide, versus dB/mm for the above-described, hypothetical non-slotted structures within the best etching and coating quality we could achieve. Using these three ingredients, one can conclude an improvement of times compared to a non-finger-loaded waveguide. The metrics are estimated to be still at least twice higher for the aforementioned slot waveguide thanks to smaller , at the cost of the described challenges.
@@
The algorithms used for this study are standard and are outlined in “Methods.” The corresponding authors can provide code scripts upon reasonable request.
@@ Acknowledgements
The Natural Sciences and Engineering Research Council of Canada (NSERC), the SiEPICfab consortium, the B.C. Knowledge Development Fund (BCKDF), the Canada Foundation for Innovation (CFI), MITACS funding with Dream Photonics Inc., and the Schmidt Sciences funded this research.
@@
This work was conceptualized by I.T., L.C., and S.S. C.P., J.S., J.Y., and N.J. contributed materials and technical information. At the same time, M.H. helped with design and S.J., K.M.A., D.W., and I.T. built the organic Si phase shifters described in this paper. I.T. conducted measurements while N.J. and O.E. helped with the modulator measurement setup. The manuscript was written by I.T., L.C., and S.S., with assistance from all authors.